Operations Report – February 15th

Crew IKEA Operations Report 15 Feb 2018

SOL: 5

Name of person filing report: Shannon Rupert

Non-nominal systems: Power System Control
Notes on non-nominal systems: Nothing new to report

Generator (hours run): Generator was once again run all day

Solar— not checked % SOC
Diesel – 72 %
Propane – 47 psi.
Ethanol Free Gasoline (5 Gallon containers for ATV) – 0 gallons.

Water (trailer) – 0 gallons.
Water (static) – 425 gallons
Trailer to Static Pump used – no
Water (loft) – Static to Loft Pump used – yes
Water Meter: not recorded

Toilet tank emptied: no

ATV’s Used: no
Oil Added? n/a
ATV Fuel Used: n/a Gals
# Hours the ATVs were Used today: 0
Notes on ATVs: Nothing to report

Deimos rover used: Yes
Hours:
Beginning charge: 100
Ending charge: 96
Currently charging: Yes

Sojourner rover used: No
Hours:
Beginning charge:
Ending charge:
Currently charging: yes

Spirit rover used: Yes
Hours: 20.2
Beginning charge: 100
Ending charge: 75
Currently charging: No

Opportunity rover used: Yes
Hours: 14.9
Beginning charge:
Ending charge:
Currently charging: Yes

Curiosity rover used: No
Hours:
Beginning charge:
Ending charge:
Currently charging: No

HabCar used and why, where? Yes, to town for packages. New jumper cables arrived for HabCar. Other set is in Hab.

General notes and comments: Septic tank is almost full. Pumping it out is scheduled for Monday. New battery charger (to replace the one that arrived broken) arrived today and will be stored in the RAM.

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: The front airlock door was left unsecured while the contractors were working and the window got blown out again. We will fix it before the crew arrives.

The new stairs are taking shape! Both of the ladders have been installed on the upper deck and the base of the new stairs is in place but not installed. The ladder sections and the lower stairs still need to be installed as well as the railing at the top of the stairs. The contractor expects the installation to be complete on Saturday but the polishing/painting will be done later.

I should be able to send photos tomorrow.

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Electrical connection has been scheduled for 3-6 weeks out.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing at this time

Operations Report – February 14th

Crew IKEA Operations Report 14 Feb 2018

SOL: 4

Name of person filing report: Shannon Rupert

Non-nominal systems: Power System Control

Notes on non-nominal systems: Nothing new to report

Generator (hours run): Due to clouds and contractors, the generator was run all day. The guys needed power for welding and sanding, so even when the sun did come out, we decided to just keep it on.

Solar— 99 % SOC

Diesel – 72 %

Propane – 48 psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – 0 gallons.

Water (trailer) – 0 gallons.

Water (static) – 450 gallons (added 80 gallons)

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: not recorded

Toilet tank emptied: no

ATV’s Used: yes

Oil Added? n/a

ATV Fuel Used: n/a Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: I started and warmed them up today.

Deimos rover used: Yes

Hours: moved to another parking space

Beginning charge:

Ending charge:

Currently charging: No

Sojourner rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used: Yes

Hours: Moved to another parking space

Beginning charge:

Ending charge:

Currently charging: No

Opportunity rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: Yes

Curiosity rover used: Yes

Hours: Moved to another parking space

Beginning charge:

Ending charge:

Currently charging: No

HabCar used and why, where? Yes, moved to Hab to pump the remaining water into the static tank. My truck is dead (apparently the battery has gone bad as there is battery acid on top) and so tomorrow we will use the HabCar to go to town and get a new battery for the truck and also pick up the packages that are in town.

General notes and comments: Still muddy most of the day but a strong wind later in the afternoon has dried things out quickly.

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: The new stairs and ladders arrived! The contractors began installation and one of the two ladders has been put into place. They are currently bare metal only. We will need to decide which looks best, polished steel or paint. It looks like we will not have time to finish them before the crew arrives on Saturday, so we won’t have to decide right away. Finishing them will probably happen in early May.

Summary of GreenHab operations: Atila continued working on the aquaponics set up and the reorganization of plants. We harvested some greens for salads tonight, and eyed the cucumbers. They may not last until the crew arrives! In addition, we doubled up the pots for the plants in an effort to maintain moisture in the soil longer. In the future, potting soil should be mixed with perlite to ameliorate this problem.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing at this time

Operations Report – February 13th

Crew IKEA Operations Report 13 Feb 2018

SOL: 3

Name of person filing report: Shannon Rupert

Non-nominal systems: Power System Control

Notes on non-nominal systems: nothing new to report

Generator (hours run): Turned off at 3:30 pm, on at 5:55 pm

Solar— SOC not known %

Diesel – did not check %

Propane – did not check psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – 0 gallons.

Water (trailer) – 0 gallons.

Water (static) – 350 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – no

Water Meter: not recorded

Toilet tank emptied: no

ATV’s Used: none

Oil Added? n/a

ATV Fuel Used: n/a Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: nothing to report

Deimos rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: No

Sojourner rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: Yes

Opportunity rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: Yes

Curiosity rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: No

HabCar used and why, where? No

General notes and comments: It rained overnight and was cloudy most of
the day, which was why we kept the generator on most of the day and
did not check some things. Here’s hoping the mud will dry tomorrow.
Everything comes to a standstill in the mud. We did get much work
done on the IKEA project.

Summary of internet: nothing to report

Summary of suits and radios: nothing to report

Summary of Hab operations: nothing to report

Summary of GreenHab operations: nothing to report

Summary of ScienceDome operations: nothing to report

Summary of RAMM operations: nothing to report

Summary of health and safety issues: When it is muddy here, it is very
hard to walk. I successfully did go outside and did not fall but I
did not want to go into the buildings and take all the mud in with me.

Questions, concerns and requests to Mission Support: None at this time.

Operations Report – February 12th

Crew IKEA Operations Report 12 Feb 2018

SOL: 2

Name of person filing report: Shannon Rupert

Non-nominal systems: Power System Control

Notes on non-nominal systems: nothing new to report

Generator (hours run): Turned off at 9:30 am, on at 5:00 pm

Solar— SOC 82 %

Diesel – 75 %

Propane – 49 psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – 0 gallons.

Water (trailer) – 0 gallons.

Water (static) – 350 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: not recorded

Toilet tank emptied: no

ATV’s Used: none

Oil Added? n/a

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: Nothing to report

Deimos rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: No

Sojourner rover used: Yes

Hours:

Beginning charge: 100

Ending charge: 73

Currently charging: yes

Spirit rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: Yes

Opportunity rover used: Yes

Hours: 14.3

Beginning charge: 100

Ending charge: 64

Currently charging: Yes

Curiosity rover used: Yes

Hours: 12.5

Beginning charge: 100

Ending charge: 98

Currently charging: No

HabCar used and why, where? Yes, to town for water

General notes and comments: We continued working on the Ikea project today.

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Atila continues his work on the GreenHab and will be sending a GreenHab Report tonight. The grow lights were operating on a different schedule than they should have been on, so we are changing the extra hours on the lights from 7-9 pm to reflect the longer days.

Summary of ScienceDome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Okay, mathletes, this is for you:

In the Hab, the kitchen has 111 inches of cabinets and they take up 130 inches of wall space. They are 24 inches deep. The Hab is 30 feet in diameter. Downstairs we want to add 240 inches of cabinets that are also 24 inches deep. How much wall space do we need?

If you are still with me, try this one. The dome is 25 feet in diameter. If we want to put 24 inch deep cabinets on ½ of the wall space, a) how much wall space is that and b) how many inches of cabinetry would we be able to fit in that space?

Extra credit to anyone who can tell us how to calculate countertop dimensions for the kitchen and lab.

We would like our numbers checked. Thank you.

Operations Report – February 11th

Crew IKEA Operations Report 11 Feb 2018

SOL: 1

Name of person filing report: Shannon Rupert

Non-nominal systems: Power System Control

Notes on non-nominal systems: Waiting for contractor to troubleshoot

Generator (hours run): Turned off at 9:15 am, on at 6:15 pm

Solar— SOC 92 %

Diesel – did not check %

Propane – did not check psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – 0 gallons.

Water (trailer) – 0 gallons.

Water (static) – 400 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – no

Water Meter: not recorded

Toilet tank emptied: no

ATV’s Used: none

Oil Added? n/a

ATV Fuel Used: n/a Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: nothing to report

Deimos rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: not sure

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging: no

Spirit rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: not sure

Opportunity rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: yes, electrical power was restored to the parking
spot on the east side of the GreenHab

Curiosity rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging: Not sure

HabCar used and why, where? No

General notes and comments: Crew 188 left the Hab at 8:30 am this
morning. They did a nice job of cleaning. We are expecting rain in
the next couple of days so will make another trip for water tomorrow.

Summary of internet: nothing to report

Summary of suits and radios: nothing to report

Summary of Hab operations: The Hab is in very good condition. We
spent much of the day working on the Ikea project and have come up
with some exciting changes that will make the kitchen and EVA areas
much easier to use and maintain. We changed the air filter in the
furnace. We are expecting the new stairs, railing and upstairs
ladders to be installed this week.

Summary of GreenHab operations: Atila spent the afternoon working on
the GreenHab. Most of the large plants were dried out and wilting
while others were overwatered. The greens were well tended. We are
limiting the tomato plants to 10, pepper plants to 5 and removing the
melon plants. This is needed in order to have room for the aquaponics
system. There is a cucumber I have my eye on, and if Atila doesn’t
pick it, we will have a nice salad in a couple of days.

Summary of ScienceDome operations: There was new damage to the floor
while Crew 188 was here. There are several small pits and one larger
one in the floor.

Summary of RAM operations: Nothing to report. We are hoping the
electrician can come soon to connect the power.

Summary of health and safety issues: nothing to report

Questions, concerns and requests to Mission Support: None at this time.

Operations Report – February 9th

Crew 188 Operational Report 09/02/2018

Sol 12

Zac Trolley

Non-Nominal Systems: electrical system, electrical cords, propane, EVA Suits

Notes on non-nominal systems: See individual systems at the end of the report.

Generator (hours run): 14.5h

Solar: SOC 80%

Diesel: 72%

Propane: 55

Ethanol Free Gasoline (5 Gallon containers for ATV): 3 Gallons

Water (trailer): 0 Gallons

Water (static): 65 Gallons

Trailer to Static Pump used: No

Water (loft) – Static to Loft Pump used: Yes

Meter: 130949.8

Toilet tank emptied: No

ATVs Used: Yes

Oil Added: No

ATV Fuel Used: 0 Gallons

# Hours the ATVs were Used today: 0.5

Notes on ATVs: The Gasoline storage area needs to be moved. The sun is creating a large pressure differential on the jerry cans, making that area an explosion risk.

Deimos rover used: No

Hours: Hours: 108.1

Beginning charge:

Ending charge:

Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: No

Hours: 19.5h

Beginning charge:

Ending charge:

Currently charging: No

Opportunity rover used: No

Hours: 19.1h

Beginning charge:

Ending charge:

Currently charging: No

Curiosity rover used: No

Hours: 12.0h

Beginning charge:

Ending charge:

Currently charging: No

HabCar: No

General notes and comments:

During an EVA, a plastic handle from the passenger side roll cage on Demos came off. We have the handle, but no hardware to re-attach it.

Generator was turned off at 9:25

Generator was turned on at 18:24

I’ve found two extension cords that have been damaged and need to be replaced. If there are 3 prong plugs available, I can repair the cords. Please advise.

Summary of internet: Tests were run on the internet connectivity and passed on to Bernard.

Summary of suits and radios: I’ve passed on a report about the EVA suits

Summary of Hab operations: Propane was smelt again this afternoon around 4pm. The heater had been turned off for 24 hours, and nothing else was on. I strongly suggest a licensed contractor investigate the propane leak before another team moves into Hab. There is a significant safety risk as we have not been able to pinpoint the source of the leak.

Summary of EVA operations: Nominal

Summary of GreenHab operations: The temperature gauge display has some non-functional LCD areas, a new one has been ordered.

Summary of Science Dome operations: The science dome is awaiting a portable heater as per the MDRS Director.

Summary of RAM operations: Not Operational.

Summary of health and safety issues: Nominal

Questions, concerns and requests to Mission Support: We could use another bag of toilet deodorizer, as we only have one puck left.

Operations Report – February 4th

Crew 188 Operational Report 04/02/2018

Sol 7

Zac Trolley

Non-Nominal Systems: electrical system, electrical cords

Notes on non-nominal systems: See individual systems at the end of the report.

Generator (hours run): 16.9h

Solar: SOC 80%

Diesel: 75%

Propane: 61 psi

Ethanol Free Gasoline (5 Gallon containers for ATV): 8 Gallons

Water (trailer): 0 Gallons

Water (static): 170 Gallons

Trailer to Static Pump used: No

Water (loft) – Static to Loft Pump used: Yes

Meter 130739.8

Toilet tank emptied: No

ATVs Used: None

Oil Added: None

ATV Fuel Used: None

# Hours the ATVs were Used today: 0

Notes on ATVs: none

Deimos rover used: No

Hours: Hours: 108.1

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: No

Hours: 18.0h

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Opportunity rover used: No

Hours: 13.5

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Curiosity rover used: No

Hours: 11.3h

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

HabCar: Not used

General notes and comments: We are rotating the Rovers with the one plug in we have, and that should be sufficient if we are taking 2 rovers at a time per EVA.

It would seem that the electrical generation isn’t being regulated that well. The diesel and solar power systems will vary between 120V and 140V. No known reason why this is happening.

During an EVA, a plastic handle from the passenger side roll cage on Demos came off. We have the handle, but no hardware to re-attach it.

Generator was turned off at 10:36

Generator was turned on at 17:28

I’ve found two extension cords that have been damaged and need to be replaced. If there are 3 prong plugs available, I can repair the cords. Please advise.

Summary of internet: Internet connectivity continues to be a problem even with the data purchased. Devices are having troubles connecting and this is limiting the crew’s ability to monitor the CapCom emails

Summary of suits and radios: The clasp on Suit 1 was bent back to place, it’s as good as new.

Suit 2 was repaired, it was a cold solder joint, as well as the connection not being crimped. I expertly re-attached the connector and the suit is operational

Summary of Hab operations: Nominal

Summary of EVA operations: In the warm weather the fogging issue isn’t as pronounced. We are hoping for warm weather to maintain visual accuracy in the helmets.

Summary of GreenHab operations: The temperature gauge display has some non-functional LCD areas, a new one has been ordered.

Summary of Science Dome operations: The science dome is awaiting a portable heater as per the MDRS Director.

Summary of RAM operations: Not Operational.

Summary of health and safety issues: Nominal

Questions, concerns and requests to Mission Support: We could use another bag of toilet deodorizer, as we only have one pack left.

Operations Report – February 7th

Crew 188 Operational Report 07/02/2018

Sol 10

Zac Trolley

Non-Nominal Systems: electrical system, electrical cords

Notes on non-nominal systems: See individual systems at the end of the report.

Generator (hours run): 15.2h

Solar: SOC 78%

Diesel: 74%

Propane: 55 – Reading taken at 2:30 PM. There are no units on the gauge.

Ethanol Free Gasoline (5 Gallon containers for ATV): 3 Gallons

Water (trailer): 0 Gallons

Water (static): 120 Gallons

Trailer to Static Pump used: No

Water (loft) – Static to Loft Pump used: Yes

Meter 130875.1

Toilet tank emptied: No

ATVs Used: None

Oil Added: Yes, about 600ml total.

ATV Fuel Used: 0 Gallons

# Hours the ATVs were Used today: 0

Notes on ATVs: The Gasoline storage area needs to be moved. The sun is creating a large pressure differential on the jerry cans, making that area an explosion risk.

Deimos rover used: Yes

Hours: Hours: 108.1

Beginning charge: 100%

Ending charge: 100%

Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: No

Hours: 18.0h

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Opportunity rover used: Yes

Hours: 19.1h

Beginning charge: 100%

Ending charge: 99%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 11.6h

Beginning charge: 100%

Ending charge: 99%

Currently charging: No

HabCar: No

General notes and comments:

During an EVA, a plastic handle from the passenger side roll cage on Demos came off. We have the handle, but no hardware to re-attach it.

Generator was turned off at 9:18

Generator was turned on at 18:20

I’ve found two extension cords that have been damaged and need to be replaced. If there are 3 prong plugs available, I can repair the cords. Please advise.

The propane smell was back today. Around 5:00pm there was two people in the Hab, and the smell was detected in the same place as before. The downstairs heater was off. It has been my understanding that the downstairs heater turns off automatically when it reaches a certain temperature. The source of the gas could be a portion of propane that is not combusted when the heater automatically turns off.

If this is NOT the way the heater functions, please let me know.

We used the detector and there was a positive detection of gas. Unfortunately, we did not think to do a control check with the instrument when no gas was smelt to ensure we are not getting false positives. I did another bubble test on all the visible gas connections, and nothing was found.

Summary of internet: Tests were run on the internet connectivity and passed on to Bernard.

Summary of suits and radios: Nominal

Summary of Hab operations: We are running low of vegetables, and there is a concern that our vegetarians will not have enough food for the remainder of the mission.

Summary of EVA operations: Nominal

Summary of GreenHab operations: The temperature gauge display has some non-functional LCD areas, a new one has been ordered.

Summary of Science Dome operations: The science dome is awaiting a portable heater as per the MDRS Director.

Summary of RAM operations: Not Operational.

Summary of health and safety issues: Nominal

Questions, concerns and requests to Mission Support: We could use another bag of toilet deodorizer, as we only have one puck left.

Operations Report – February 6th

Crew 188 Operational Report 06/02/2018

Sol 9

Zac Trolley

Non-Nominal Systems: electrical system, electrical cords

Notes on non-nominal systems: See individual systems at the end of the report.

Generator (hours run): 15.2h

Solar: SOC 74%

Diesel: 75%

Propane: 61 psi

Ethanol Free Gasoline (5 Gallon containers for ATV): 3 Gallons

Water (trailer): 0 Gallons

Water (static): 170 Gallons

Trailer to Static Pump used: No

Water (loft) – Static to Loft Pump used: Yes

Meter 130825.4

Toilet tank emptied: Yes

ATVs Used: The three blue ATVs were used today

Oil Added: None

ATV Fuel Used: 2 Gallons

# Hours the ATVs were Used today: 2

Notes on ATVs: It is very difficult to see the fuel level in the ATVs with the suit helmet on. Also, the gas can that had a little fuel in the bottom imploded due to temperature change. There was no leak, but the structural integrity of the can has been compromised and it should be destroyed. I suggest a new location be found to keep the cans out of the sun.

Deimos rover used: No

Hours: Hours: 108.1

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: No

Hours: 18.0h

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Opportunity rover used: Yes

Hours: 19.0h

Beginning charge: 100%

Ending charge: 97%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 11.4h

Beginning charge: 100%

Ending charge: 99%

Currently charging: No

HabCar: Yes, to get a load of water.

General notes and comments:

During an EVA, a plastic handle from the passenger side roll cage on Demos came off. We have the handle, but no hardware to re-attach it.

Generator was turned off at 9:00

Generator was turned on at 18:04

I’ve found two extension cords that have been damaged and need to be replaced. If there are 3 prong plugs available, I can repair the cords. Please advise.

Summary of internet: Tests were run on the internet connectivity and passed on to Bernard.

Summary of suits and radios: Nominal

Summary of Hab operations: We are running low of vegetables, and there is a concern that our vegetarians will not have enough food for the remainder of the mission.

Summary of EVA operations: Nominal

Summary of GreenHab operations: The temperature gauge display has some non-functional LCD areas, a new one has been ordered.

Summary of Science Dome operations: The science dome is awaiting a portable heater as per the MDRS Director.

Summary of RAM operations: Not Operational.

Summary of health and safety issues: Nominal

Questions, concerns and requests to Mission Support: We could use another bag of toilet deodorizer, as we only have one puck left.

The “tunnel” had a few sharp edges of the metal at the corners, and a jacket was torn. I’ve taped over the edges so it doesn’t happen again.

Operations Report – February 5th

Crew 188 Operational Report 05/02/2018

Sol 8

Crew reporting: Zac Trolley

Non-Nominal Systems: electrical system, electrical cords

Notes on non-nominal systems: See individual systems at the end of the report.

Generator (hours run): 16.9h

Solar: SOC 84%

Diesel: 75%

Propane: 61 psi

Ethanol Free Gasoline (5 Gallon containers for ATV): 5 Gallons

Water (trailer): 0 Gallons

Water (static): 150 Gallons

Trailer to Static Pump used: No

Water (loft) – Static to Loft Pump used: Yes

Meter 130788.5

Toilet tank emptied: Yes

ATVs Used: Yes, all four were used today.

Oil Added: None

ATV Fuel Used: 3 Gallons

# Hours the ATVs were Used today: 1

Notes on ATVs: It is very difficult to see the fuel level in the ATVs with the suit helmet on.

Deimos rover used: No

Hours: Hours: 108.1

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: No

Hours: 18.0h

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Opportunity rover used: No

Hours: 13.5

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Curiosity rover used: No

Hours: 11.3h

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

HabCar: Not used

General notes and comments:

During an EVA, a plastic handle from the passenger side roll cage on Demos came off. We have the handle, but no hardware to re-attach it.

Generator was turned off at 10:36

Generator was turned on at 17:49

I’ve found two extension cords that have been damaged and need to be replaced. If there are 3 prong plugs available, I can repair the cords. Please advise.

Summary of internet: Sub Nominal

Summary of suits and radios: Nominal

Summary of Hab operations: Nominal

Summary of EVA operations: Nominal

Summary of GreenHab operations: The temperature gauge display has some non-functional LCD areas, a new one has been ordered.

Summary of Science Dome operations: The science dome is awaiting a portable heater as per the MDRS Director.

Summary of RAM operations: Not Operational.

Summary of health and safety issues: Nominal

Questions, concerns and requests to Mission Support: We could use another bag of toilet deodorizer, as we only have one puck left.